OpenStudy (anonymous):

Just to make sure this is right.... Suppose f(x)=2^x. What is f(0)? Is it 1?

OpenStudy (unklerhaukus):

\[f(x)=2^x\]\[f(0)=2^0=1\] yes\[x^0=1\]

OpenStudy (anonymous):

Okay thank you!